OpenStudy (anonymous):

can someone help me with 4 and 5?

OpenStudy (anonymous):

what do you know about matrix multiplication?

OpenStudy (anonymous):

that you add the first and last numbers

OpenStudy (anonymous):

do you understand a dot product between 2 matrix?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok then, we're multiplying a 2x2 matrix by a 2x2 matrix we know that multiplication is possible because the middle are both 2 2x (2 = 2) x2 and the resulting matrix will be a 2x2 matrix ex. (1x3)*(3x1) is possible and the resulting matrix will be a 1x1 matrix for dot product, if we have 2 vectors (1x n) or (n x1) the dot product is equal to the entries multiplied to their respective entry then added together ex. \[\left(\begin{matrix}1 \\ 2 \\3 \end{matrix}\right) * \left(\begin{matrix}1 \\ 2\\3 \end{matrix}\right)\] the dot product of those 2 matrices =1*1 +2*2+3*3=14

OpenStudy (anonymous):

for matrix multiplication, if you're finding the entry (i,j) of a matrix the entry(i,j) is equal to the dot product of row i of the first matrix and column j ex.\\[A=\left[\begin{matrix}1 & 2 \\ 3 & 4\end{matrix}\right]\ B=\left[\begin{matrix}2 & 3 \\ 4 & 5\end{matrix}\right]\] for A*B the entry of (1,1) would be =1*2+2*4=10 entry of (1,2) =1*3+2*5=13 \[A*B=\left[\begin{matrix}10 & 13 \\ ? & ?\end{matrix}\right]\] and so on and so forth

OpenStudy (anonymous):

any questions?

OpenStudy (anonymous):

would 4 be a?

OpenStudy (anonymous):

5=c?

OpenStudy (anonymous):

4 is not a, just look at the first entry the dot product [1 6] and [1 -3] is not equal to 1

OpenStudy (anonymous):

5 is incorrect as well, again just look at the entry (1,1)

OpenStudy (anonymous):

4=b?

OpenStudy (anonymous):

ok before you continue guessing, what is the dot product of [1 6] and [1 -3] ?